OpenStudy (anonymous):

Use differentials to approximate the quantity. Square root of 5

OpenStudy (helder_edwin):

use this \[ \large f(x+h)=h\cdot f'(x)+f(x) \] choose x and h appropiatedly!!

OpenStudy (helder_edwin):

in your case \[ \large f(x)=\sqrt{x} \]

OpenStudy (anonymous):

It also says in the question to (Give your answer correct to 4 decimal places.). I tried 2.2361but it was incorrect.

OpenStudy (helder_edwin):

let x=4 and h=1. plug these in what i just wrote tell me what u get

OpenStudy (anonymous):

got 2.25 and put in 2.2500 and it was correct, thanks for the help

OpenStudy (helder_edwin):

u r welcome